OpenStudy (anonymous):

solve the equation... 3^(6+3x)=(1/27) the answer is -3 show me why and how please.

OpenStudy (vijay):

\[3^{6+3x} = 3^{-3}\]

OpenStudy (vijay):

6 + 3x = -3 -3x = 9 x = -3

OpenStudy (anonymous):

take logbase3 of both sides 6+3x = -3 -3x = 9 x = -3